1.       A woman launches a boat from one shore of a straight river and wants to land at the point directly on the opposite shore. If the speed of the boat (relative to the water) is 10 mi/h and the river is flowing east at the rate of 5 mi/h, in what direction should she head the boat in order to arrive at the desired landing point?




1
Expert's answer
2022-03-14T20:05:10-0400

For the woman wanted at a point directly across the river. She needs to move in a horizontal direction on component O.

5+10cosθ=O

Cos θ=-5/10

Cosθ=-1/2

θ =1200

Hence, she should head the boat in the direction of θ =1200